OpenStudy (marsmcnugget):

What is the value of b? 11+4b=8b-13 B= ?

OpenStudy (zzr0ck3r):

subtract 4b from both sides add 13 to both sides divide both sides by 4

OpenStudy (anonymous):

4b-8b=-13-11 -4b=-24 b=6

OpenStudy (marsmcnugget):

thanks